A copper cable is 1.2km long and has a cross-sectional Area of 5mm².Find it's resistance at 80°C if at 20°C the resistivity of copper is 0.02×10^-6 ohm m and it's temperature coefficient of resistance is 0.004/°C.


1
Expert's answer
2021-07-21T09:52:12-0400

Resistance at 20°C "R_{20}=\frac{\rho{l}}{a} = \frac{0.02×10^{-6}×1.2×10^3M}{5×10{^-6}}=4.8ohms"

Resistance at 80°C "R_{80}=R_{20}[1+\alpha_{20}(80-60)] = 4.8[1+0.004(60)]= 5.95Ω"

"R_{80}= 5.95\u03a9"